OpenStudy (anonymous):

in a classroom there are two kinds of seats: single and double. There are twice as many single seats as double seats. The room can accommodate 76 pupils. How many of each kind of seats are there?

OpenStudy (blacksteel):

Let us suppose there are x single seats and y double seats. We know that x = 2y, and we know that x + 2y = 76. Using these 2 equations, we can solve for x and y: x + 2y = 76 2y + 2y = 76 4y = 76 y = 19 x = 2(19) = 38. So there are 19 double seats and 38 single seats.

OpenStudy (anonymous):

thanks :]